You borrow $1,000 from the bank and agree to repay the loan over the next year in 12 equal monthly payments of $90. However, the bank also charges you a loan initiation fee of $28, which is taken out of the initial proceeds of the loan. What is the effective annual interest rate on the loan, taking account of the impact of the initiation fee? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places. Use a financial calculator or Excel.)

WHAT IS Effective annual interest rate %???

plz solve this Q
Answer #1

Hi

Here Borrowing amount = $1,000

loan initiation fees = $28

So Borrowing amount after initiation fees PV= 1000 - 28 = $972

time nper = 1 years = 12 months

monthly payment PMT = $90

So using excel function RATE we will find monthly rate

Excel formulas below

=RATE(C3,-C4,C2,,) 1 972 112 190 nper PMT monthly rate Effective Annual Rate =RATE(C3,-C4,C2,) (1+C5}^12-1

Excel Values below

=RATE(C3,-C4,C2,,) PV nper PMT monthly rate Effective Annual Rate $972 12 $90 1.66% 21.83%

So effective annual rate is 21.83%

Thanks